# FACTR: Force-Attending Curriculum Training for Contact-Rich Policy Learning

### v1

@@ Abstract
Many contact-rich tasks humans perform, such as box pickup or rolling dough, rely on force feedback for reliable execution. However, this force information, which is readily available in most robot arms, is not commonly used in teleoperation and policy learning. Consequently, robot behavior is often limited to quasi-static kinematic tasks that do not require intricate force-feedback. In this paper, we first present a low-cost, intuitive, bilateral teleoperation setup that relays external forces of the follower arm back to the teacher arm, facilitating data collection for complex, contact-rich tasks. We then introduce FACTR, a policy learning method that employs a curriculum which corrupts the visual input with decreasing intensity throughout training. The curriculum prevents our transformer-based policy from over-fitting to the visual input and guides the policy to properly attend to the force modality. We demonstrate that by fully utilizing the force information, our method significantly improves generalization to unseen objects by 43% compared to baseline approaches without a curriculum. Video results and instructions at https://jasonjzliu.com/factr/
@@ I Introduction
To mitigate this imbalance, we propose Force-Aware Curriculum Training (FACTR), a curriculum training strategy designed to improve the policy’s ability to effectively leverage force information. FACTR systematically reduces the reliance on visual information during training by applying operators such as Gaussian blurring or downsampling with varying scales to visual inputs. A scheduler gradually decreases the blurring scale and increases the fidelity of the visual inputs. Intuitively, this approach encourages the policy to focus more on force input during initial training phases and gradually balances force with visual inputs as training progresses. We ground this intuition with a theoretical analysis on a simplified scenario through the framework of Neural Tangent Kernels [11]. We explore FACTR in both the pixel space and latent space, testing various operators and scheduling strategies. Our experiments show that FACTR improves the success rate for unseen objects by an average of 40.0% in four challenging contact-rich tasks (Fig. LABEL:fig:teaser Right) – box lifting, prehensile pivoting, fruit pick-and-place, and dough rolling – showcasing the efficacy of our force-attending curriculum training.
@@ III-E Overall Control Law for the Leader Arm
IV FACTR: Force-Aware Curriculum Training
@@
To fully leverage the robot force data collected from our teleoperation system, we introduce Force-Aware Curriculum Training (FACTR), a training strategy designed to effectively integrate force information into policy learning. FACTR applies operators like Gaussian blur or downsampling to corrupt visual information, where the amount of visual corruption decreases throughout training. The curriculum intuitively encourages contribution from the force modality at the start of training. We ground this intuition with a theoretical analysis on a simplified scenario through the framework of Neural Tangent Kernels [11].
@@ IV-A Problem Statement and Base Model
IV-B Force-Aware Curriculum
@@ V-C Policy Evaluation
  • ACT (Vision+Force): Action Chunking Transformer which takes in both visual and force observation, but trained without a curriculum.
@@
For each object in each task, we evaluated 5-10 trials. We present the average success rate for training and testing objects, respectively. Detailed evaluation results for each object can be found in the Appendix XI.
@@
FACTR leads to better recovery behavior. Another notable observation is that FACTR also facilitates recovery behavior. Specifically, we evaluate the box-lifting task with five trials per object. A trial begins when the policy successfully lifts the box for the first time; we then knock the box down and assess the second attempt. As shown in Table I, all policies maintain nearly 100% recovery success on training objects. However, for test objects, the vision-only policy’s success rate drops significantly from 31.7% on the first attempt to 13.3% on the second. In contrast, force-aware policies maintain similar success rates across both attempts.
@@ VIII Cost Analysis of Our Teleoperation System with Force Feedback
Please see Table III for a detailed Bill of Materials and breakdown of the cost to create one leader arm and leader gripper as part of our teleoperation with force feedback system. This is accurate pricing as of the paper submission.
@@
Dynamixel XM430-W210-T 2 \$540
Dynamixel XC330-T288-T 6 \$540
U2D2 Control PCB 1 \$20
12V 20A Power Supply 1 \$25
PLA Printer Plastic N/A \$10
Total \$1135
TABLE III: We present the bill of materials of one leader arm teleoperation device with force feedback. The total cost is around \$1135.
@@ X Behavior Cloning Policy Architecture and Training Hyperparameters
XI Detailed Quantitative Results
We present the detailed evaluation results for each task in TABLE V, VI, VII, and VIII.
@@ XI Detailed Quantitative Results
TABLE V: Comparison of methods for Box Lift task.
@@
TABLE VI: Comparison of methods for Non-Prehensile Pivot task.
@@
TABLE VII: Comparison of methods for Fruit Pick-Place task.
@@
TABLE VIII: Comparison of methods for Rolling Dough task.

### v2

@@ Abstract
Many contact-rich tasks humans perform, such as box pickup or rolling dough, rely on force feedback for reliable execution. However, this force information, which is readily available in most robot arms, is not commonly used in teleoperation and policy learning. Consequently, robot behavior is often limited to quasi-static kinematic tasks that do not require intricate force-feedback. In this paper, we first present a low-cost, intuitive, bilateral teleoperation setup that relays external forces of the follower arm back to the teacher arm, facilitating data collection for complex, contact-rich tasks. We then introduce FACTR, a policy learning method that employs a curriculum which corrupts the visual input with decreasing intensity throughout training. The curriculum prevents our transformer-based policy from over-fitting to the visual input and guides the policy to properly attend to the force modality. We demonstrate that by fully utilizing the force information, our method significantly improves generalization to unseen objects by 43% compared to baseline approaches without a curriculum. Video results, codebases, and instructions at https://jasonjzliu.com/factr/
@@ I Introduction
To mitigate this imbalance, we propose Force-Attending Curriculum Training (FACTR), a curriculum training strategy designed to improve the policy’s ability to effectively leverage force information. FACTR systematically reduces the reliance on visual information during training by applying operators such as Gaussian blurring or downsampling with varying scales to visual inputs. A scheduler gradually decreases the blurring scale and increases the fidelity of the visual inputs. Intuitively, this approach encourages the policy to focus more on force input during initial training phases and gradually balances force with visual inputs as training progresses. We ground this intuition with a theoretical analysis on a simplified scenario through the framework of Neural Tangent Kernels [11]. We explore FACTR in both the pixel space and latent space, testing various operators and scheduling strategies. Our experiments show that FACTR improves the success rate for unseen objects by an average of 40.0% in four challenging contact-rich tasks (Fig. LABEL:fig:teaser Right) – box lifting, prehensile pivoting, fruit pick-and-place, and dough rolling – showcasing the efficacy of our force-attending curriculum training.
@@ III-E Overall Control Law for the Leader Arm
IV FACTR: Force-Attending Curriculum Training
@@
To fully leverage the robot force data collected from our teleoperation system, we introduce Force-Attending Curriculum Training (FACTR), a training strategy designed to effectively integrate force information into policy learning. FACTR applies operators like Gaussian blur or downsampling to corrupt visual information, where the amount of visual corruption decreases throughout training. The curriculum intuitively encourages contribution from the force modality at the start of training. We ground this intuition with a theoretical analysis on a simplified scenario through the framework of Neural Tangent Kernels [11].
@@ IV-A Problem Statement and Base Model
IV-B Force-Attending Curriculum
@@ V-C Policy Evaluation
  • ACT (Vision+Force) [13, 14]: Action Chunking Transformer which takes in both visual and force observation, but trained without a curriculum.
@@
For each object in each task, we evaluated 5-10 trials. We present the average success rate for training and testing objects, respectively. Detailed evaluation results for each object can be found in the Appendix XII.
@@
FACTR leads to better recovery behavior. Another notable observation is that FACTR also facilitates recovery behavior. Specifically, we evaluate the box-lifting task with five trials per object. A trial begins when the policy successfully lifts the box for the first time; we then knock the box down and assess the second attempt. As shown in Table I, all policies maintain nearly 100% recovery success on training objects. However, for test objects, the vision-only policy’s success rate drops significantly from 31.7% on the first attempt to 13.3% on the second. In contrast, force-attending policies maintain similar success rates across both attempts.
@@ VIII Cost Analysis of Our Teleoperation System with Force Feedback
Please see Table III for a detailed Bill of Materials and breakdown of the cost to create one leader arm and leader gripper as part of our teleoperation with force feedback system. This is accurate pricing as of the paper release.
@@
Dynamixel XM430-W210-T 2 \$539.80
Dynamixel XC330-T288-T 6 \$539.40
U2D2 Control PCB 1 \$32.10
12V 20A Power Supply 1 \$24.99
FPX330-S102 Servo Bracket 1 \$8.70
Polymaker PLA PRO Filament 1 \$24.99
U2D2 Power Hub Board 1 \$19.99
14AWG Cable 1 \$23.99
3/4” Bearing 1 \$6.99
Screws - \$15.99
Total \$1229.95
TABLE III: We present the bill of materials of one leader arm teleoperation device with force feedback. The total cost is around \$1229.95.
@@ X Behavior Cloning Policy Architecture and Training Hyperparameters
XI Additional Experiments
Adaptive Layer Norm as an Alternative to a Curriculum As an alternative to a curriculum, we experimented with Adaptive Normalization Layers [23] on the action tokens conditioned on the force input to improve the force conditioning (AdaNorm in Tab. V). However, we found that it creates instability in training and leads to overfitting.
Data Augmentation as alternative to curriculum. We run 10 data augmentation experiments varying noise probabilities and noise levels applied to vision input. The best of these policies, shown as DataAug in Tab. V. We find that policies trained with lower augmentation levels perform well on train objects but worse than FACTR on test objects, likely because low levels of augmentations fail to facilitate proper attention to force. Higher levels of augmentation policies do not perform well across train and test sets, likely because they fail to leverage vision altogether. Our curriculum approach eliminates such tradeoff between vision and force, allowing the policy to attend properly to both force and vision.
FACTR AdaNorm NoiseAug
Train (%) 90.0 25.0 85.0
Test (%) 77.7 6.1 65.0
TABLE V: Pivot task training and testing performance.
More test objects. We doubled the number of testing objects used in three of our tasks, visualized in Fig. 8 and updated the success rate comparison figure in Tab. VI. By conducting testing on more unseen objects and comparing with the best performing baseline (Bi-ACT), we observe that our method yields an improved generalization performance.
Box Lift Pivot Rolling Dough
ACT (Vision-Only) 35/120 30/130 0/60
Bi-ACT 68/120 76/130 41/60
FACTR (Ours) 105/120 101/130 46/60
TABLE VI: Policy evaluation on unseen objects across 3 tasks.
Figure 8: We more than doubled the testing set size across 3 tasks.
XII Detailed Quantitative Results
We present the detailed evaluation results for each task in TABLE VII, VIII, IX, and X.
@@ XII Detailed Quantitative Results
TABLE VII: Comparison of methods for Box Lift task.
@@
TABLE VIII: Comparison of methods for Non-Prehensile Pivot task.
@@
TABLE IX: Comparison of methods for Fruit Pick-Place task.
@@
TABLE X: Comparison of methods for Rolling Dough task.